# What Do Privacy Advertisements Communicate to Consumers?

### v1

@@ Abstract.
When companies release marketing materials aimed at promoting their privacy practices or highlighting specific privacy features, what do they actually communicate to consumers? In this paper, we explore the impact of privacy marketing materials on: (1) consumers’ attitude towards the organizations providing the campaigns, (2) overall privacy awareness, and (3) the actionability of suggested privacy advice . To this end, we investigated the impact of four privacy advertising videos and one privacy game published by five different technology companies. We conducted 24 semi-structured interviews with participants randomly assigned to view one or two of the videos or play the game. Our findings suggest that awareness of privacy features can contribute to positive perceptions of a company or its products. The ads we tested were more successful in communicating the advertised privacy features than the game we tested. We observed that advertising a single privacy feature using a single metaphor in a short ad increased awareness of the advertised feature. The game failed to communicate privacy features or motivate study participants to use the features. Our results also suggest that privacy campaigns can be useful for raising awareness about privacy features and improving brand image, but may not be the most effective way to teach viewers how to use privacy features.
@@
journalyear: 2024journalvolume: 4journalnumber: 2024doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
In recent years, an increasing number of companies have used their privacy practices as a way of differentiating themselves from competitors. Apple launched an advertisement focusing on the iPhone “Ask App Not to Track” feature in May 2022 (O’Flaherty, 2022). Experts estimate this new iPhone feature caused a \$12 billion dollar decline in revenue for Meta (O’Flaherty, 2022). The advertisement shows the protagonist, Ellie, stumbling upon an auction where her information is up for sale. As Ellie toggles various privacy protection features on her iPhone, bidders begin to disappear until she is the only one left in the auction room (see Figure 1, and Figure 10 in Appendix A.3 for illustration) (Apple, 2022). Apple is not alone in showcasing their privacy protection features as a way to gain consumer trust; Samsung, another smartphone giant, launched an advertisement with similar message in the same year (Samsung, 2022). Google and WhatsApp also have ads highlighting their security and privacy features (Google Ads, 2020; WhatsApp, 2022), and Twitter released an old-school video game in an attempt to help users gain better understanding of their privacy settings (Twitter, 2022).
While experts have praised some of these privacy campaigns (O’Flaherty, 2022; Ulanoff, 2022) and criticized others (Hurlerand and Wodinsky, 2022; Winkie, 2022), it is unclear what consumers take away from these campaigns. In this study we explore what consumers exposed to corporate privacy campaigns are learning from these campaigns, and what impacts these campaigns have on consumers’ privacy awareness and attitudes.
@@
Our exploratory study suggests that awareness of privacy features can contribute to positive perceptions of a company and its products. The ads we tested were more effective than the game in communicating privacy features to our participants. We observed that advertising a single privacy feature using a single metaphor in a short ad increased awareness of the advertised feature among our participants. The game performed poorly in both explaining privacy features and encouraging user engagement with the features. Our results suggest that privacy campaigns might not be the most effective method for educating users on how to use privacy features. In addition, despite their positive impressions, most non-users in our study were reluctant to switch products after seeing ads due to concerns about switching costs, data synchronization, and familiarity with their current operating system. Nonetheless, over time, such campaigns may lead to more interest in a brand’s products, and might be particularly effective for products and services with low switching costs.
@@ 2.1. Effective Advertisements
Video advertisements contain scenes and imagery carefully crafted by advertisers. Mohanty and Ratneshwar examined the factors that influence a viewer’s subjective comprehension of visual metaphors in graphic ads (Mohanty and Ratneshwar, 2015). The researchers found that incongruity, which is the discrepancy between what is actually shown and what it is referring to, could lead to poor comprehension (Mohanty and Ratneshwar, 2015). Mohanty and Ratneshwar also conducted a later study investigating the effects of incongruity on ad effectiveness, and found that a moderate level of incongruity results in the most effective ads, recommending that advertisers be cautious when selecting visual metaphors (Mohanty and Ratneshwar, 2016).
Some video ads are narrative ads that contain a story or plot. Kim et al. found that narrative ads were effective because they generated emotive responses, offered high entertainment value, and created credibility (Kim et al., 2017). Laurence conducted a study researching the role of storytelling in narrative ads and found that when used to induce positive emotions, participants had more positive brand attitudes (Laurence, 2018). Manyiwa and Ross looked at the impact of negative emotions in advertising and found that participants with higher reported self-efficacy had more favourable views towards ads that used fear (Manyiwa and Brennan, 2012).
Privacy marketing has the potential to reach a large audience but prior work has not examined its effectiveness. Our research seeks to gain insights into consumer privacy awareness by examining recent privacy marketing campaigns and what consumers taking away from them. We investigated how the privacy campaign impacted viewers’ impression of the organization and how well participants understood privacy metaphors.
2.2. Privacy and Security Games
@@
Video games can be an effective way of communicating a simplified, simulated model of any concept to a player (Persuasive Games, 2023). To educate users about phishing, Sheng et al. developed “Anti-Phishing Phil,” and found games to be an effective educational method (Sheng et al., 2007). The researchers credited the effectiveness of the treatment to the interactive game content (Sheng et al., 2007). Maqsood et al. investigated whether web based games could help children aged 11 to 13 practice safer online behaviors. They found significant improvements in knowledge and intended behaviors after being exposed to the game and found that the game was “usable, fun, and relatable” (Maqsood et al., 2018). Google also launched the “Interland” game aimed at educating children about online safety, along with YouTube videos and other online resources to help parents and teachers raise these issues with children (Moscaritolo, 2017). Thompson and Irvine evaluated “CyberCiege,” a security game that teaches students basic cybersecurity concepts and found that students did not treat the game as a lab session, often overlooking the lab manual provided and instead played straight away. Researchers noted this shift in mentality is something they would keep in mind for future renditions of “CyberCiege” (Thompson and Irvine, 2011).
@@ 2.3. Privacy Awareness
Major themes that emerged from the selected campaigns were social media privacy settings, online privacy and behavioral advertising, and end-to-end encryption in instant messaging. In this section, we explore privacy awareness research related to each theme.
@@
Acquisti and Gross conducted one of the earliest studies about privacy awareness and information sharing practices on Facebook and found that while most participants claimed to be aware of their profile visibility, there were “a significant minority of members [who] are unaware of those tools and options” (Acquisti and Gross, 2006). Another Facebook privacy study, conducted by Sohoraye et al., focused on users’ knowledge of privacy statements and privacy regulations in their countries. The researchers found that while many people are concerned about their personal information, they are relatively unaware of related privacy information (Sohoraye et al., 2015). Users’ lack of awareness of social media privacy setting could lead to misaligned expectation between what they intended to share and reality (Liu et al., 2011; Madejski et al., 2012).
@@
Common social media privacy concerns include personal safety, reputation management, social disclosure and how easy it may be for other users to see their activities (Ahern et al., 2007). Depending on the platform and its intended audience, users have different privacy concerns. Jeong and Kim found that Facebook users are more concerned about what others may post on their timeline, versus Twitter users who are more concerned about the content of tweets they have posted (Jeong and Kim, 2017). Baruh et al. found that participants’ privacy concern did not significantly correlate with their actual social media usage (Baruh et al., 2017). Building on top of Baruh et al’s findings, Barth et al. investigated the privacy paradox and found that users prioritized functionality and usability above possible privacy concerns, regardless of their technical ability or financial situation (Barth et al., 2019).
@@
Increased privacy awareness may also lead to increased concerns about data used for behavioural advertising. Researchers have found that while users are generally aware of tracking, their understanding of how personalized advertising works may not reflect reality (Rader, 2014). Personalized ads and behavioral advertising is considered as both a useful feature, but also one that may bring privacy risks (Ur et al., 2012) or become a nuisance (Dehling et al., 2019).
@@
When it comes to instant messaging, security and privacy may have minor influence on why users choose a particular app. DeLuca et al. found that overall, peer influence had the biggest impact on user choice in messaging apps (De Luca et al., 2016). Abu-Salma et al. interviewed 60 participants to investigate barriers in secure communication tool adoption and found that most participants lacked understanding about end-to-end encryption, potentially limiting their motivation to adopt secure messaging tools (Abu-Salma et al., 2017). Other barriers to adoption could include misconceptions that SMS or email may be more secure (Stransky et al., 2021), or lack of trust in end-to-end encryption (Dechand et al., 2019). Stransky et al. investigated how visualization may impact perception of messaging security and found that perception of security depended more on the reputation of the app itself (Stransky et al., 2021). When attempting to improve user understanding of end-to-end encryption, Bai et al. found that it is most effective to use simple wording with minimal technical detail (Bai et al., 2020).
@@
Privacy advertisements and campaigns are designed to increase privacy awareness, which can impact consumer choices. Kelley et al. examined privacy as part of a user’s decision to download or install an app. The researchers found that users seldom considered privacy information when making a decision to download apps. However, when the authors introduced privacy information into the app store, users were much more likely to consider privacy as a factor in their decision making (Kelley et al., 2013). Tsai et al. found that when privacy information is prominently and intuitively visible, people would be willing to pay at a premium for privacy (Tsai et al., 2011). Similarly, Emami-Naeini et al., found that consumers are willing to pay more for Internet-of-Things products with better privacy and security practices, such as de-identified cloud storage (Emami-Naeini et al., 2023).
@@ 2.4. Acceptance of Privacy and Security Advice
Redmiles et al. developed a set of quality metrics for security and privacy advice: perceived actionability, perceived efficacy, and comprehensibility. The researchers found that generally, while all study participants found the advice actionable, neither the users nor the experts were particularly good at prioritizing advice (Redmiles et al., 2020). Research in security awareness campaigns also helps to identify key factors that may influence the success or failure of a campaign. Bada et al. evaluated security campaigns across the UK and Africa. The researchers found that security education “needs to be targeted, actionable, doable and provide feedback,” and that invoking fear is an ineffective way of promoting security awareness (Bada et al., 2015). Building on Bada et al.’s work, Das et al. conducted a comprehensive literature review on research related to security and privacy awareness. Based on their review, the researchers developed the Security and Privacy Acceptance Framework (SPAF), identifying three key barriers that could impact a user’s acceptance or rejection of recommended security and privacy advice, being awareness, motivation and ability (Das et al., 2022).
Technology has also affected the way in which people learn new information. Kross et al. surveyed a nationally representative sample of US adults, and found that online safety, security, and privacy are 3 of the 19 categories of subjects people tried to learn more about through online resources. The researchers identified YouTube as the most commonly used learning channel (Kross et al., 2021). Akgul et al. built on Kross et al.’s work and analyzed influencer videos on YouTube containing VPN ads. They found that “VPN ads likely reach billions of viewers, comparable to ambitious industry efforts at influencing users’ understandings of security and privacy tools.” However, their analysis of these ads suggested that the information they contain may not always be correct and may potentially have negative impacts on viewers’ understanding of online safety (Akgul et al., 2022).
@@ 3. Methods
We conducted an exploratory study, aimed at investigating the impact of privacy marketing materials on privacy perception. Specifically, we explored how privacy marketing materials, such as privacy ads and games, impact viewers’ attitude towards the organizations providing the campaign, awareness of privacy features and issues, and the actionability of suggested privacy advice. In the subsections below, we first introduce the privacy marketing campaigns that were selected for this study, then outline the recruitment and screening procedure, as well as the interview protocol and data analysis method.
@@ 3.1. Privacy Marketing Materials
We searched for privacy marketing materials on YouTube and the web using the keywords “Privacy Advertisements,” “Privacy Ads,” and “Privacy Awareness.” We found five campaigns that were published in 2022 by technology companies. We selected one video ad from each of four ad campaigns we found. We also selected the Twitter DataDash game, which had received some media attention in 2022. Here we provide a short summary of each campaign. Table 1 provides more information about each campaign. Additional images from the campaigns are presented in Appendix A.3.
@@
Protagonist stumbles upon a room packed with people, where an auctioneer is selling off her private information to the highest bidder, shown in Figure 1. As the information for sale becomes more sensitive, she decides to take action. By using Apple’s “Ask App Not To Track,” she eliminates half of the bidders in the room. Next, by using Apple’s “Protect Mail Activity,” she empties the room completely, including all remaining bidders, the auction attendants, as well as the auctioneer (Apple, 2022).
Figure 1. Auctioneer facilitating auction of protagonist’s data in Apple ad.
@@
Protagonist knows that much of her life is on her Samsung phone, where she struggles to get away from floating balloons representing apps, shown in Figure 2. Utilizing Samsung’s privacy dashboard, she takes control of her privacy. By using Samsung’s permission manager, she turns off precise location, illustrated with a yellow umbrella from the Weather app, to blend into a sea of yellow umbrellas. As she continues using her phone, she chooses to grant or deny location and microphone permission for apps as needed. Using Samsung’s privacy dashboard, she is able to review all app permission usage in a central location. Samsung’s Knox Security is mentioned at the end of the video (Samsung, 2022).
Figure 2. Protagonist struggles to get away from balloons representing apps in Samsung ad.
@@
People arrive at the post office wanting to mail their letters. To their disbelief, the clerk, shown in  Figure 3, insists on sending those letters using a messenger pigeon, assuring his customers that using pigeons as a delivery method is “mostly” secure. The ad states that 5.5 billion texts per day are sent without encryption, while with WhatsApp, their conversations would not be one of them (WhatsApp, 2022).
Figure 3. Post office clerk (right) uses carrier pigeon for sending letters in WhatsApp ad.
@@
Singer in a Google T-shirt singing an altered version of the song “Every Breath You Take” (The Police, 2023a, b), where lyrics were changed to be relevant to web tracking (e.g. “Every click you make, I will be watching you.”) As the singer shoulder-surfs people while they browse the Internet, they are visually uncomfortable with the singer’s presence, as shown in Figure 4. One person decides to start using DuckDuckGo, causing the singer to be dragged away by an invisible force. An announcer explains that “The Internet does not have to be so creepy” and introduces DuckDuckGo products (DuckDuckGo, 2022).
Figure 4. Singer wearing Google T-shirt (right) shoulder-surfs a person browsing the Internet in DuckDuckGo ad.
@@
DataDash is an online game simulating the style of 90s old-school platform games, where players journey through PrivaCity with Data dog as illustrated in Figure 5. Players must collect five artifacts to pass to the next level. The artifacts for level 1 and 2 are envelopes containing dog bones. At the end of each level, players are shown a short explanation referring to a particular privacy and safety feature, with a “Twitter Setting” button that takes them to a specific privacy and safety setting on Twitter. The privacy and safety settings for each level are: (1) Ad Preference, (2) Direct Messages, (3) Location Information, and (4) Audience and Tagging . The website where DataDash is hosted explains the purpose of the game: “The Twitterverse can be tricky to navigate if you don’t know your way around. So we made a game to help you understand our privacy policy a little better” (Twitter, 2022).
Figure 5. Start of level 1 of Twitter game.
@@
Apple Data Auction Published on YouTube May 18, 2022 1 min 35 sec https://www.youtube.com/watch?v=NOXK4EVFmJY
Samsung You’re in Control Published on YouTube Feb 9, 2022 1 min 35 sec https://www.youtube.com/watch?v=kZKK80urUZc
WhatsApp A New Era of Personal Privacy with Default End-to-End Encryption Published on YouTube Oct 17, 2022 1 min https://www.youtube.com/watch?v=zvI4cVGWJhM
DuckDuckGo Watching You Published on YouTube May 17, 2022 30 sec https://www.youtube.com/watch?v=QWpPyYlZXNI
Twitter DataDash Estimated to be May, 2022 based on time of media coverage (Winkie, 2022; Hurlerand and Wodinsky, 2022). Not Applicable https://twitterdatadash.com

### v2

@@ Abstract.
When companies release marketing materials aimed at promoting their privacy practices or highlighting specific privacy features, what do they actually communicate to consumers? In this paper, we explore the impact of privacy marketing on: (1) consumers’ attitudes toward the organizations providing the campaigns, (2) overall privacy awareness, and (3) the actionability of suggested privacy advice . To this end, we investigated the impact of four privacy advertising videos and one privacy game published by five different technology companies. We conducted 24 semi-structured interviews with participants randomly assigned to view one or two of the videos or play the game. Our findings suggest that awareness of privacy features can contribute to positive perceptions of a company or its products. The ads we tested were more successful in communicating the advertised privacy features than the game we tested. We observed that advertising a single privacy feature using a single metaphor in a short ad increased awareness of the advertised feature. The game failed to communicate privacy features or motivate study participants to use the features. Our results also suggest that privacy campaigns can be useful for raising awareness about privacy features and improving brand image, but may not be the most effective way to teach viewers how to use privacy features.
@@
journalyear: 2024journalvolume: 2024journalnumber: 4doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
In recent years, an increasing number of companies have used their privacy practices as a way of differentiating themselves from competitors. Apple launched a video advertisement focusing on the iPhone “Ask App Not to Track” feature in May 2022. Experts estimate this new iPhone feature caused a 12 billion dollar decline in revenue for Meta (O’Flaherty, 2022). The advertisement shows the protagonist, Ellie, stumbling upon an auction where her information is up for sale. As Ellie toggles various privacy protection features on her iPhone, bidders begin to disappear until she is the only one left in the auction room (see Figure 1, and Figure 9 in Appendix A.3 for illustration) (Apple, 2022). Apple is not alone in showcasing their privacy protection features as a way to gain consumer trust; Samsung, another smartphone giant, launched an advertisement with a similar message in the same year (Samsung, 2022). Google and WhatsApp also have ads highlighting their security and privacy features (Google Ads, 2020; WhatsApp, 2022), and Twitter released an old-school video game in an attempt to help users gain a better understanding of their privacy policies (Damien, 2022; Twitter, 2022).
While experts have praised some of these privacy ad campaigns (O’Flaherty, 2022; Ulanoff, 2022) and criticized others (Hurlerand and Wodinsky, 2022; Winkie, 2022), it is unclear what consumers take away from these campaigns. In this study, we explore what consumers exposed to corporate privacy ad campaigns are learning from these campaigns, and what impacts these campaigns have on consumers’ privacy awareness and attitudes.
@@
Our exploratory study suggests that awareness of privacy features can contribute to positive perceptions of a company and its products. The ads we tested were more effective than the game in communicating privacy features to our participants. We observed that advertising a single privacy feature using a single metaphor in a short ad increased awareness of the advertised feature among our participants. The game performed poorly in both explaining privacy features and encouraging user engagement with the features. Our results suggest that privacy campaigns might not be the most effective method for educating users on how to use privacy features. In addition, despite their positive impressions, most non-users in our study were reluctant to switch products after seeing ads due to concerns about switching costs, data synchronization, and familiarity with their current operating system. Nonetheless, over time, such campaigns may lead to more interest in a brand’s products and might be particularly effective for products and services with low switching costs.
@@ 2.1. Effective Advertisements
Video advertisements contain scenes and imagery carefully crafted by advertisers. Mohanty and Ratneshwar examined the factors that influence a viewer’s subjective comprehension of visual metaphors in graphic ads. The researchers found that incongruity, which is the discrepancy between what is actually shown and what it is referring to, could lead to poor comprehension (Mohanty and Ratneshwar, 2015). Mohanty and Ratneshwar also conducted a later study investigating the effects of incongruity on ad effectiveness, and found that a moderate level of incongruity results in the most effective ads, recommending that advertisers be cautious when selecting visual metaphors (Mohanty and Ratneshwar, 2016).
Some video ads are narrative ads that contain a story or plot. Kim et al. found that narrative ads were effective because they generated emotive responses, offered high entertainment value, and created credibility (Kim et al., 2017). Laurence conducted a study researching the role of storytelling in narrative ads and found that when used to induce positive emotions, participants had more positive brand attitudes (Laurence, 2018). Manyiwa and Ross looked at the impact of negative emotions in advertising and found that participants with higher reported self-efficacy had more favorable views towards ads that used fear (Manyiwa and Brennan, 2012).
Privacy marketing has the potential to reach a large audience but prior work has not examined its effectiveness. Our research seeks to gain insights into consumer privacy awareness by examining recent privacy marketing campaigns and what consumers are taking away from them. We investigated how the privacy campaigns impacted viewers’ impressions of the organizations and how well participants understood privacy metaphors.
$2.$ Privacy and Security Games
@@
Video games can be an effective way of communicating a simplified, simulated model of any concept to a player (Persuasive Games, 2023). Sheng et al. developed the “Anti-Phishing Phil” game and found it to be an effective way to educate people to identify phishing URLs. The researchers credited the game’s effectiveness to its interactive content (Sheng et al., 2007). Maqsood et al. investigated whether web-based games could help children aged 11 to 13 practice safer online behaviors. They found significant improvements in knowledge and intended behaviors after being exposed to the game and found that the game was “usable, fun, and relatable” (Maqsood et al., 2018). Google also launched the “Interland” game aimed at educating children about online safety, along with YouTube videos and other online resources to help parents and teachers raise these issues with children (Moscaritolo, 2017). Thompson and Irvine evaluated “CyberCiege,” a security game that teaches students basic cybersecurity concepts. They found that students did not treat the game as a lab session, often overlooking the lab manual provided and instead playing straight away. Researchers noted this shift in mentality is something they would keep in mind for future renditions of “CyberCiege” (Thompson and Irvine, 2011).
@@ 2.3. Privacy Awareness
Major themes that emerged from the campaigns we selected for our study were social media privacy settings, online privacy and behavioral advertising, and end-to-end encryption in instant messaging. In this section, we explore privacy awareness research related to each theme.
@@
Acquisti and Gross conducted one of the earliest studies about privacy awareness and information-sharing practices on Facebook and found that while most participants claimed to be aware of their profile visibility, there were “a significant minority of members [who] are unaware of those tools and options” (Acquisti and Gross, 2006). Another Facebook privacy study, conducted by Sohoraye et al., focused on users’ knowledge of privacy statements and privacy regulations in their countries (Sohoraye et al., 2015). The researchers found that while many people are concerned about their personal information, they are relatively unaware of related privacy information (Sohoraye et al., 2015). Users’ lack of awareness of social media privacy settings could lead to misaligned expectations between what they intended to share and reality (Liu et al., 2011; Madejski et al., 2012).
@@
Common social media privacy concerns include personal safety, reputation management, social disclosure, and how easy it may be for other users to see their activities (Ahern et al., 2007). Depending on the platform and its intended audience, users have different privacy concerns. Jeong and Kim found that Facebook users have more privacy concerns related to what others may post on their timeline, versus Twitter users who are more concerned about the privacy of tweets they have posted (Jeong and Kim, 2017). Baruh et al. found that participants’ privacy concerns did not significantly correlate with their actual social media usage (Baruh et al., 2017). Building on top of Baruh et al’s findings, Barth et al. investigated the privacy paradox and found that users prioritized functionality and usability above possible privacy concerns, regardless of their technical ability or financial situation (Barth et al., 2019).
@@
Increased privacy awareness may also lead to increased concerns about data used for behavioral advertising. Researchers have found that while users are generally aware of tracking, their understanding of how personalized advertising works may not reflect reality (Rader, 2014). Personalized ads and behavioral advertising are considered useful features, but also ones that may bring privacy risks (Ur et al., 2012) or become a nuisance (Dehling et al., 2019).
@@
When it comes to instant messaging, security and privacy may have a minor influence on why users choose a particular app. DeLuca et al. found that overall, peer influence had the biggest impact on user choice in messaging apps (De Luca et al., 2016). Abu-Salma et al. interviewed 60 participants to investigate barriers to secure communication tool adoption and found that most participants lacked understanding about end-to-end encryption, potentially limiting their motivation to adopt secure messaging tools (Abu-Salma et al., 2017). Other barriers to adoption could include misconceptions that SMS or email may be more secure (Stransky et al., 2021), or lack of trust in end-to-end encryption (Dechand et al., 2019). Stransky et al. investigated how visualization may impact the perception of messaging security and found that the perception of security depended more on the reputation of the app itself (Stransky et al., 2021). When attempting to improve user understanding of end-to-end encryption, Bai et al. found that it is most effective to use simple wording with minimal technical detail (Bai et al., 2020).
@@
Privacy advertising campaigns are designed to increase privacy awareness, which can impact consumer choices. Kelley et al. examined privacy as part of a user’s decision to download or install an app. The researchers found that users seldom considered privacy information when making a decision to download apps. However, when the authors introduced privacy information into the app store, users were much more likely to consider privacy as a factor in their decision-making process (Kelley et al., 2013). Tsai et al. found that when privacy information was prominently visible, people were willing to pay a premium for privacy (Tsai et al., 2011). Similarly, Emami-Naeini et al. found that consumers were willing to pay more for Internet-of-Things products with better privacy and security practices, such as de-identified cloud storage (Emami-Naeini et al., 2023).
@@ 2.4. Acceptance of Privacy and Security Advice
Redmiles et al. developed a set of quality metrics for security and privacy advice: perceived actionability, perceived efficacy, and comprehensibility. The researchers found that generally, while all study participants found the advice actionable, neither the users nor the experts were particularly good at prioritizing advice (Redmiles et al., 2020). Research on security awareness campaigns also helps to identify key factors that may influence the success or failure of a campaign. Bada et al. evaluated security campaigns across the UK and Africa. The researchers found that security education “needs to be targeted, actionable, doable and provide feedback,” and that invoking fear is an ineffective way of promoting security awareness (Bada et al., 2015). Subsequently, Das et al. conducted a comprehensive literature review on research related to security and privacy awareness. Based on their review, the researchers developed the Security and Privacy Acceptance Framework (SPAF), identifying three key barriers that could impact a user’s acceptance or rejection of recommended security and privacy advice: awareness, motivation, and ability (Das et al., 2022).
Technology has also affected the way in which people learn new information. Kross et al. surveyed a nationally representative sample of US adults and found that online safety, security, and privacy are three of the 19 categories of subjects people tried to learn more about through online resources. The researchers identified YouTube as the most commonly used learning channel (Kross et al., 2021). Akgul et al. analyzed influencer videos on YouTube containing VPN ads. They found that “VPN ads likely reach billions of viewers, comparable to ambitious industry efforts at influencing users’ understandings of security and privacy tools.” However, their analysis of these ads suggested that the information they contain may not always be correct and may potentially have negative impacts on viewers’ understanding of online safety (Akgul et al., 2022).
@@ 3. Methods
We conducted an exploratory interview study, aimed at investigating the impact of privacy marketing materials on privacy perception. Specifically, we explored how privacy marketing materials, such as privacy ads and games, impact viewers’ attitudes towards the organizations providing the campaign, awareness of privacy features and issues, and the actionability of suggested privacy advice. In the subsections below, we first introduce the privacy marketing campaigns that were selected for this study, then outline the recruitment and screening procedure, as well as the interview protocol and data analysis method.
@@ 3.1. Privacy Marketing Materials
We searched for privacy marketing materials on YouTube and the web using the keywords “Privacy Advertisements,” “Privacy Ads,” and “Privacy Awareness.” We found four campaigns that were published in 2022 by technology companies. We selected one video ad from each of the four ad campaigns we found. We also selected the Twitter DataDash game, which received some media attention in 2022. Here we provide a short summary of each campaign. Table 1 provides more information about each campaign. Additional images from the campaigns are presented in Appendix A.3.
@@
The protagonist stumbles upon a room packed with people where an auctioneer is selling off her private information to the highest bidder, as shown in Figure 1. As the information for sale becomes more sensitive, she decides to take action. Using Apple’s “Ask App Not To Track,” she eliminates half of the bidders in the room. Next, by using Apple’s “Protect Mail Activity,” she empties the room, including all remaining bidders, the auction attendants, as well as the auctioneer (Apple, 2022).
Figure 1. The auctioneer facilitating the auction of the protagonist’s data in the Apple ad.
@@
The protagonist knows that much of her life is on her Samsung phone, where she struggles to get away from floating balloons representing apps, as shown in Figure 2. Utilizing Samsung’s privacy dashboard, she takes control of her privacy. By using Samsung’s permission manager, she turns off the precise location, illustrated with a yellow umbrella from the Weather app, to blend into a sea of yellow umbrellas. As she continues using her phone, she chooses to grant or deny location and microphone permission for apps as needed. Using Samsung’s privacy dashboard, she is able to review all app permission usage in a central location. Samsung’s Knox Security is mentioned at the end of the video (Samsung, 2022).
Figure 2. The protagonist struggles to get away from balloons representing apps in the Samsung ad.
@@
People arrive at the post office wanting to mail their letters. To their disbelief, the clerk, as shown in Figure 3, insists on sending those letters using a messenger pigeon, assuring his customers that using pigeons as a delivery method is “mostly” secure. The ad states that 5.5 billion texts per day are sent without encryption, while with WhatsApp, their conversations would not be one of them (WhatsApp, 2022).
Figure 3. The post office clerk (right) uses a carrier pigeon for sending letters in the WhatsApp ad.
@@
A singer in a Google T-shirt singing an altered version of the song “Every Breath You Take” (The Police, 2023), where lyrics were changed to be relevant to web tracking, e.g., “Every click you take … I will be watching you.” As the singer shoulder-surfs people while they browse the Internet, they are visually uncomfortable with the singer’s presence, as shown in Figure 4. One person decides to start using DuckDuckGo, causing the singer to be dragged away by an invisible force. An announcer explains that “The Internet does not have to be so creepy” and introduces DuckDuckGo products (DuckDuckGo, 2022).
Figure 4. The singer wearing a Google T-shirt (right) shoulder-surfs a person browsing the Internet in the DuckDuckGo ad.
@@
DataDash is an online game simulating the style of 90s old-school platform games, where players journey through PrivaCity with data dog as illustrated in Figure 5. Players must collect five artifacts to pass to the next level. The artifacts for levels 1 and 2 look like envelopes containing dog bones. At the end of each level, players are shown a short explanation referring to a particular privacy and safety feature, with a “Twitter Settings” button that takes them to a specific privacy and safety setting on Twitter. The privacy and safety settings for each level are: (1) Ad Preference, (2) Direct Messages, (3) Location Information, and (4) Audience and Tagging . The website where DataDash is hosted explains the purpose of the game: “The Twitterverse can be tricky to navigate if you don’t know your way around. So we made a game to help you understand our privacy policy a little better” (Twitter, 2022).
Figure 5. The start of level 1 of the Twitter game.
@@
Apple Data Auction Apple’s YouTube channel on May 18, 2022 1 min, 35 sec https://www.youtube.com/watch?v=NOXK4EVFmJY
Samsung You’re in Control Samsung’s YouTube channel on Feb 9, 2022 1 min, 35 sec https://www.youtube.com/watch?v=kZKK80urUZc1
WhatsApp A New Era of Personal Privacy with Default End-to-End Encryption WhatsApp’s YouTube channel on Oct 17, 2022 1 min https://www.youtube.com/watch?v=zvI4cVGWJhM2
DuckDuckGo Watching You DuckDuckGo’s YouTube channel on May 17, 2022 30 sec https://www.youtube.com/watch?v=QWpPyYlZXNI
Twitter DataDash A blog post by Twitter’s Privacy Center (Damien, 2022) on May 11, 2022 Not Applicable https://twitterdatadash.com